OpenStudy (anonymous):

What is the value of h in the function f(x)=1/2(x=5)2-9? A. 5 B. 1/2 C. 9 D. -5

OpenStudy (across):

I am very confused: Why do you have a second equals sign in the middle of the function, and where is h?

OpenStudy (anonymous):

they didnt give me an h value :-/ and that was supposed to be x=5

OpenStudy (across):

I am now more confused. n_n

myininaya (myininaya):

what? I'm confused too. The way you typed your problem is that exactly the way it looks?

OpenStudy (anonymous):

yeah. but the 2 is small and its raised

myininaya (myininaya):

and inside the parenthesis is x=5?

myininaya (myininaya):

\[f(x)=\frac{1}{2}(x \pm 5)^2-9\] i think it should either be + or -

OpenStudy (across):

\[f(x)=\frac{1}{2}(x=5)^2-9...\]I can't seem to make any sense of this.

myininaya (myininaya):

and I'm assuming you want to find the x coordinate of the parabola

OpenStudy (anonymous):

Ummm, I guess.. I dont know. Im really really bad at math

myininaya (myininaya):

can you write the function down one more time?

myininaya (myininaya):

and i meant x coordinate of the vertex of the parabola lol

OpenStudy (anonymous):

f(x)=1/2(x+5)2-9

myininaya (myininaya):

ok great!

myininaya (myininaya):

now it makes more sense

OpenStudy (anonymous):

oh. sorry! haha

myininaya (myininaya):

\[f(x)=\frac{1}{2}(x+5)^2-9\] \[f(x)=a(x-h)^2+k\] what is h?

myininaya (myininaya):

well we want -h=5 right?

myininaya (myininaya):

so if -h=5, then h=?

OpenStudy (anonymous):

-5? lol sorry, im super weak in math

myininaya (myininaya):

-5 is tight!

myininaya (myininaya):

i mean right*

OpenStudy (anonymous):

YAY!! :D

OpenStudy (anonymous):

can you help me with more? Ive got a ton

myininaya (myininaya):

Ok

myininaya (myininaya):

But you have to give me a sprite or dr. pepper

OpenStudy (anonymous):

I'll give you both!

myininaya (myininaya):

lol

myininaya (myininaya):

ok thanks for the imaginary soda